OpenStudy (anonymous):

th e area of a square is 121 square meters. What is the length of one side of this square? A. 12.1 m B. 14,641 m C. 60.5 m D. 11 m

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (anonymous):

Dan did a great job explaining, did you not understand?

OpenStudy (here_to_help15):

D

OpenStudy (here_to_help15):

Its really really simple

OpenStudy (here_to_help15):

11*11 is.......

OpenStudy (here_to_help15):

or 121/11

OpenStudy (here_to_help15):

@shelbydancer

OpenStudy (mathstudent55):

The area of a square is the length of the side squared. \(A = s^2\) In this case, \(121 ~m^2 = s^2\) \(s^2\) means s multiplied by itself. \(s \times s = 121\) What number multiplied by itself equals 121?